Jim has designed a rectangle with an area of 100 sq feet and a perimeter of 401 ft. What are the dimentions of Jims rectangle?
Answer by mananth(16946)

Jim has designed a rectangle with an area of 100 sq feet and a perimeter of 401 ft. What are the dimentions of Jims rectangle?
let the length be x
and the width be y.
2x+2y = 401
and xy = 100
y= 100/x
2x+ 2*100/x=401
2x^2 +200 = 401x
2x^2-401x+200=0
2x^2-400x -x +200=0
2x(x-200)-1(x-200)=0
(x-200)(2x-1)=0
x= 200 or 1/2

The rectangle will have dimensions of 200 by 1/2